OpenStudy (anonymous):

Using De Moivre's theorem, find (-1-i)^3

OpenStudy (anonymous):

\[(-1-i)^3=(\sqrt{2}e^{-(3/4)pi})^3\]

OpenStudy (anonymous):

But my textbook says -8 ...

OpenStudy (anonymous):

but (-1-i)^3 = 2-2i

OpenStudy (anonymous):

Using my calculator, I got (-1-i)^3 expanded as: \[-3i ^{2}\sqrt{3}+9i ^{2}-3i \sqrt{3}+1\] You cancel off the first and third expression, leaving behind 9i^2 and +1 this will be -8 I want to get this by using De Moivre's theorem

OpenStudy (anonymous):

(-1-i)^3 = -(1+3i+3i^2+i^3) = -1-3i+3+i i don't know why −3i^2√3+9i^2−3i√3+1